OpenStudy (anonymous):

The difference of two numbers is 4. 5 times the smaller number is four more than three times the larger. Find the two numbers

OpenStudy (anonymous):

If one number is x, then the other number will be (4 + x) 5x = 4 + 3(x + 4) 5x = 4 + 3x + 16 Solve for x here...

OpenStudy (anonymous):

If the following equations reflect the problem statement, then @waterineyes solution is incorrect. Let x and y be the larger and the smaller number respectively. Solve the following for x and y:\[\{x-y=4,5y=4+3x\}\]\[\{x=12,y=8\} \]

OpenStudy (anonymous):

@radar How is it going "Old Timer" ?

OpenStudy (radar):

@robtobey, going well, we are experiencing a "heatwave" will be 106 tomorrow. Hope things are better in CA.

OpenStudy (radar):

I see the mistake that @waterineyes made, they did not distribute the 3 properly in 3(x+4)

OpenStudy (anonymous):

That is what I hear on TV News. My wife and I were at Bad Water, Death Valley, California, some years ago when it was 119.5 as measured on a Yellow Back low distillation thermometer ( a real thermometer). Death Valley holds the record for the highest reliably reported temperature in the Western hemisphere, 134 °F (56.7 °C) at Furnace Creek on July 10, 1913, just short of the world record, 136 °F (57.8 °C) in 'Aziziya, Libya, on September 13, 1922. Bad Water is 19 miles South of Furnace Creek.

OpenStudy (radar):

Remind me not to take a vacation in the summer at Furnace Creek.

OpenStudy (anonymous):

Maybe not so good in the Summer, however, if you have an opportunity to visit Death Valley, the geography is absolutely spectacular. Thanks you for the medal radar. Have a good day.

OpenStudy (radar):

You're welcome, it was the best answer. I spent a few weeks at Camp Irwin, CA I did get to see Death Valley, but fortunately did not spend much time in that valley.

OpenStudy (radar):

Good day to you.

OpenStudy (anonymous):

Sorry @radar, Ya there is mistake in multiplying that is 3*4 = 12 So solution is correct @robtobey. 5x = 4 + 3x + 12 2x = 16 x = 8 and (x + 4) = 12.....